Shushma has a limited income to spend on orange and apple. Market price of orange (Po) is
Rs.20 per Kg. and price of apple (Pa) is 25 per Kg. When she draws her budget line, it
terminates X axis (where orange is plotted) at 50 units. Find Shushma’s income and slope
of her budget line.
1
Expert's answer
2021-09-03T12:05:03-0400

The intercept of x-axis determines the quantity of good that can be purchased if all the income is spent on good represented on the x- axis. 

The point at which budget line terminates on x-axis is the intercept of x-axis.

Intercept of x - axis is calculated as follows:

"Intercept \space of\space x − axis=\frac {Income}{Price} \\50=\frac{ Income}{20} \\Income= \$1000"

The income of consumer is $1000.

The slope of budget line is calculated as follows:

"Slope \space of \space budget\space line =−\frac{ Price \space of \space oranges}{Price\space of\space apples}\\=−\frac{20}{25}\\=−0.8"

Slope of budget line is -0.8.